A rectangle is twice as long as it is wide. If its length is decreased by 4 and its width is decreased by 2, its area is decreased by 32. Find the original dimensions.
Answer by orca(409)

Suppose its width is x, then its length is 2x.
So its area is 2x^2.
The new rectangle's width is x + 2, length is 2x + 4, so its area is (x+2)(2x+4).
As the area of new rectangle is 32 more than the original one, we have
(x+2)(2x+4) = 2x^2 + 32
To Solve the equation for x, first we simplify it by dividing both sides by 2.
(x+2)(x+2) = x^2 + 16
Or written as
(x+2)^2 = x^2 + 16
Expanding the left side, we have
x^2 + 4x + 4 = x^2 + 16
4x = 12
x = 3
So the dimensions of the original rectangle is 3 and 6.